# Word length predicts word order: “Min-max”-ing drives language evolution

### v2

@@ 2.2 Study 2: Predictive validity of word length
Given the observation that relative lengths of nouns and verbs differ in corpora for languages with different word orders, we can ask how predictive this observation is. That is, can we predict the word order of a language based on the relative lengths of its nouns and verbs? If we are strictly interested in identifying basic word order from corpora, it is clear that the N1 ratio already allows us to do so. However, we are also interested in understanding how noun and verb lengths may be related to word order. Being able to predict word order only from noun and verb lengths in a corpus for both a modern language and its ancestral language would indicate that such properties are related to how word order evolves over time.
To investigate this question, for Study 2 I sourced additional corpora of tagged data from two pairs of historical and modern languages in which the modern language changed its word order:
@@
All languages belong to the Semitic branch of the Afroasiatic language family, a necessary constraint given the limited amount of data available for most pre-modern languages. In the case of Hebrew, the UDT has POS-tagged sentences for both Ancient [^11] https://universaldependencies.org/treebanks/hbo_ptnk and Modern [^12] https://universaldependencies.org/treebanks/he_htb varieties. In the case of Classical Arabic I used the Quranic Arabic Corpus [^13] Version 0.4: https://corpus.quran.com and for Egyptian Arabic I used the BOLT Egyptian Arabic Treebank (Maamouri et al. 2018). [^14] https://catalog.ldc.upenn.edu/LDC2018T23 (Discussion Forum) If the lengths of nouns and verbs in these corpora allows us to correctly classify Ancient Hebrew and Classical Arabic as VS and the modern varieties as SV, this would suggest that these are salient properties affecting how word order has developed diachronically. Importantly, given the textual differences between the taggedPBC (New Testament texts) and the test corpora (Biblical, Quranic, modern, and informal texts), such a finding would be significant.
In order to conduct this test, the data for Ancient Hebrew (ISO 639-3: hbo), Modern Hebrew (ISO 639-3: heb), Classical Arabic (cla) [^15] Classical Arabic does not have an assigned ISO 639-3 code or Glottocode. Accordingly, I use the abbreviation cla here, as there is no current language assigned to this code in the ISO 639-3 tables. and Egyptian Arabic (ISO 639-3: arz) were processed and the respective word classes were counted and measured as per the method outlined for Study 1 (each corpus > 400 sentences, see Table 1). The average length of nouns and verbs (including frequency effects) were then used as continuous features for a Gaussian Naive Bayes (GNB) classifier.
@@
Three additional features were also used to expand the feature dimensions. The first is a categorical feature: whether a language’s nouns are longer than its verbs. The second is a continuous feature: the difference in length between nouns and verbs. This second measure is related to minimization of complexity, based on the idea that a larger difference between lengths of word classes may effect the likelihood of that particular class being placed first. The third measure is the ratio of noun and verb lengths, and is related to maximization of complexity (via “entropy”), whereby a lower ratio indicates that the particular word classes are less likely to be easily differentiable (and therefore predictable) based on their length.
Using these five features (nounlen, verblen, nlonger, nlen-vlen, nlen/vlen), a GNB classifier was trained on the original combined word order dataset from the taggedPBC containing hand-coded classification of 1,150 languages (Modern Hebrew and Modern Arabic were removed), and then used to predict word order for our pairs of languages: Ancient Hebrew and Modern Hebrew on the one hand, Classical Arabic and Egyptian Arabic on the other. Ancient Hebrew was classified as VS while Modern Hebrew was classified as SV by the trained classifier. Classical Arabic was classified as VS while Egyptian Arabic was classified as SV.
As an additional check, I sourced data from another historical language, Old Irish, as represented in the Codex Paulinus Wirziburgensis, whereby Latin text of St. Paul’s epistles was annotated by Irish scholars during the middle of the eighth century. The dataset (Table 2; Doyle 2018) contains over 3,600 Irish glosses/sentences. As an Indo-European language, Old Irish is widely surmised to have changed its word order to verb-initial from an earlier state, so this check allows us to verify whether the same properties hold in relation to lengths of nouns and verbs. Indeed, the classifier accurately classified Old Irish as VS, based on the data extracted from the Würzburg Glosses.
@@
Study 2 thus indicates that the relative lengths of nouns and verbs, when frequency is taken into account, can predict word order from corpora unrelated to the type of corpora contained in the taggedPBC, with various implications for our understanding of language processing and language change. This is a striking finding, and I am unaware of any other research that has been able to show predictive validity of a model related to word order change. However, it is possible that this result is conditioned to some degree by the fact that the majority of the test cases are Semitic languages spoken in the same geographical area, so we now turn to an investigation of whether genealogical descent or location provides a better explanation for the correlation of noun and verb lengths with word order.
@@ 2.3 Study 3: Testing variance predicted by word length
Dunn et al. (2011) showed that descent from a common ancestor was a better predictor of many grammatical patterns than other explanations. Although Jäger & Wahle (2021) counter this claim with additional data, we can also use the current dataset to see if descent is more explanatory of word order for a given language than the lengths of nouns and verbs. Additionally, in more recent work, Guzmán Naranjo & Becker (2022) show that areal affects are as important as family membership in predicting linguistic phenomena, finding that the results in Dunn et al can be just as easily predicted by language area. This is also a claim that we can test with our dataset. Accordingly, Study 3 can proceed with several (motivated) simplifying observations.
First, for each language we can take the highest level node of family membership that is known based on current scholarship (via Glottolog). [^16] Glottolog (Hammarström et al. 2024; https://www.glottolog.org) contains expert annotations for language family affiliations for over 7,000 languages, along with additional aggregated data from a variety of sources. This is because all modern languages are the same age, and while their pathways/histories may differ, this does not impact whether they can be considered part of the same phylogenetic tree. That is, if descent is more explanatory than word class length, we would expect languages belonging to the same tree to have a higher likelihood of sharing the same word order than not. Second, we can use the macroareas identified by Glottolog as well as the actual GPS coordinates of languages to control for areal affects. Third, we can run various analyses with differing sets of data in order to see how this affects the statistical robustness of the findings. It may be, for example, that using a smaller sample gives a different result than a larger or more diverse one. Here, the more languages we have for a language family or geographical location, the higher the likelihood that any relationship identified is not due to chance. Finally, we can remove languages with “free” word order from consideration, as this leaves us with a binary distinction in word order (SV, VS) that opens up more possibilities for statistical tools.
@@
In particular, a binary dependent variable for word order means we can use hierarchical linear regression (HLR), which allows us to observe the relative impact of particular features that are theoretically relevant to the question at hand by seeing how they influence the performance of successive models. In the case of word order, we can train a regression model with a feature known to be relevant to word order (the N1 ratio), and then subsequent models that incorporate additional features of interest (“language area/location”, “family membership”, “noun/verb lengths”). We can then compare the models to see whether the additional features increase the variance explained in the data via the F-value: an increase in F-value (change) and the significance of this change indicates more/additional variance explained by the particular feature added in a model.
@@
Accordingly, for all languages in the taggedPBC I retrieved the family membership and geographical information from Glottolog, filling in missing data via Ethnologue and other sources. Then I filtered the data on four conditions: a) languages found in Dunn et al that are also present in the taggedPBC (120 languages), b) languages found in the taggedPBC that are in the four families investigated by Dunn et al (662 languages), c) languages in families with more than 75 members in the taggedPBC (1182 languages), and d) languages in families with 2 or more members in the taggedPBC (1,805 languages). For each of these conditions, I ran an HLR, adding features of interest progressively to see how they impacted the fit of the successive models and whether particular features account for additional variance.
@@
The base model (Model 1) uses the N1 ratio as the primary feature for classification. This is because we know from comparison with expert determinations of word order (see Ring 2025) that the N1 ratio significantly correlates with the classification of languages as SV or VS. Due to the strong relationship observed between language area and typological properties of language (see Guzmán Naranjo & Becker 2022), Model 2 then includes GPS coordinates and macroarea identifiers for each language as features. Model 3 additionally incorporates descent (language family membership) as a feature (as per Dunn et al. 2011), and Model 4 includes the average lengths of nouns and verbs (as per the relationship identified in this paper). If descent is more explanatory than the lengths of nouns/verbs, this should be clear from the differences between the respective models for each condition. Results are displayed in Table 3.
@@
1 (N1 ratio) 120 45.27 6.4e-10
2 (Lg area) 120 24.29 1.3e-14 12.78 2.87e-07 **
3 (Family) 120 20.02 2.7e-14 2.05 0.154
4 (N/V len) 120 15.75 2.9e-14 3.17 0.045 *
@@
1 (N1 ratio) 662 189.99 3.6e-38
2 (Lg area) 662 95.65 4.1e-64 50.07 3.7e-29 **
3 (Family) 662 78.23 2.4e-64 5.79 0.016 *
4 (N/V len) 662 65.43 2.6e-71 21.30 1.0e-09 **
@@
1 (N1 ratio) 1182 378.83 2.1e-73
2 (Lg area) 1182 210.38 3.3e-136 116.99 2.6e-66 **
3 (Family) 1182 169.09 1.3e-135 2.72 0.099
4 (N/V len) 1182 127.47 2.3e-139 14.03 9.4e-07 **
@@
1 (N1 ratio) 1805 535.41 6.1e-104
2 (Lg area) 1805 178.89 3.8e-129 46.52 5.8e-29 **
3 (Family) 1805 143.49 2.3e-128 1.66 0.197
4 (N/V len) 1805 108.16 2.2e-132 14.46 5.8e-07 **
@@
From these analyses we can see that, as expected, the N1 ratio significantly predicts word order under all conditions. Additionally, language area is a significant predictor of word order that adds to the strength of the classification. Language family (descent) does not add additional predictive power to our model under most conditions (over language area), whereas the length of nouns and verbs does. This analysis shows that the length of nouns and verbs in a language accounts for more variance and at the same time is more significant than that language’s family membership in determining its word order classification, under all test conditions. When we remove language area from the linear regression this is still the case, [^17] See additional statistical results at the linked repositories (Fn 2; Fn 5). indicating that while language area is a stronger predictor than family membership, lengths of nouns and verbs is yet a stronger one.
This largely aligns with Jäger and Wahl’s (2021) finding regarding the (lower) correlation between descent and linguistic structures in comparison to other explanations, with the caveat that Jäger and Wahl included language isolates, which is likely to greatly diminish the affect of descent. If a large number of languages in a sample are not related to others, clearly descent can have only a minimal impact on word order. By including only language families with two or more members in the taggedPBC (which excludes 111 languages), I show that descent does not explain much variance above and beyond language location/area, in contrast with noun and verb lengths. This also suggests that language family/descent is highly correlated with linguistic area, which conforms with expectations related to human history and migration. Additionally, the amount of variance explained seems to depend on the number of languages (and language families) in the sample, highlighting the need for more data for languages from under-represented language families.
@@ 3 Discussion
That arguments overall are shorter in languages of the taggedPBC, across all word orders, suggests that there are certain (non-Noun) arguments affecting these observations. Particularly, the effect of PRON seems quite likely, whereby a tendency for pronouns to be shorter crosslinguistically, along with a preference for pronouns as arguments, may impact the observation that arguments are shorter overall than predicates. This pattern aligns with the processing claim that “heavier” words (or word classes in this case) occur later in a sentence, partly validating Hypothesis 1 of study 1. The counter-argument, also from the data, is that nouns are longer than verbs in SV languages and shorter than verbs in VS languages when frequency is considered, validating Hypothesis 2 of study 1. This suggests that speakers prefer to place heavier (more “contentful”, or perhaps “novel”) constituents before lighter ones crosslinguistically in many cases.
@@
These two pressures can be (and, in fact, have been) considered as two different kinds of “efficiency”: production efficiency vs communicative efficiency. Under this view, production efficiency is related to length of words (as per Wasow 2002, 2022; Hawkins 2014) - speakers use shorter words whenever they can to efficiently produce utterances (see also Mahowald et al. 2013). The counter-pressure, communicative efficiency, is related to making a speaker’s utterances more readily accessible/interpretable by recipients (as per Levshina & Moran 2021; Levshina 2023) - speakers prioritize words that contain more meaningful, informative content in order to aid communication. Once context has been established, these longer words can be replaced by more efficient (shorter) terms such as demonstratives or pronouns. [^18] For example, there is a common crosslinguistic pattern of using a noun first to establish reference, then switching to pronouns, i.e. "that tree, it is big, it will fall…" See also discussion of the (lack of) encoding of subjects crosslinguistically in Berdicevskis et al. (2020).

### v3

@@ 2.2 Study 2: Predictive validity of word length
Given the observation that relative lengths of nouns and verbs differ in corpora for languages with different word orders, we can ask how predictive this observation is. That is, can we predict the word order of a language based on the relative lengths of its nouns and verbs? If we are strictly interested in identifying basic word order from corpora, it is clear that the N1 ratio already allows us to do so. Study 1 has shown that word length features derived from corpora are also highly significant for differentiating word orders. This is relevant because we are interested in understanding how noun and verb lengths may impact how word order changes over time. Given two sets of corpora from two timepoints of a language, being able to predict word order from noun and verb lengths in a corpus for both the modern language and its ancestral language would indicate that such properties are related to how word order evolves.
To investigate whether this is the case, for Study 2 I sourced additional corpora of tagged data from two pairs of historical and modern languages in which the the descendent changed its word order from the ancestral state:
@@
All languages belong to the Semitic branch of the Afroasiatic language family, a necessary constraint given the limited amount of data available for most pre-modern languages. In the case of Hebrew, the UDT has POS-tagged sentences for both Ancient [^14] https://universaldependencies.org/treebanks/hbo_ptnk and Modern [^15] https://universaldependencies.org/treebanks/he_htb varieties. In the case of Classical Arabic I used the Quranic Arabic Corpus [^16] Version 0.4: https://corpus.quran.com and for Egyptian Arabic I used the BOLT Egyptian Arabic Treebank (Maamouri et al. 2018), [^17] https://catalog.ldc.upenn.edu/LDC2018T23 (Discussion Forum) which are also hand-tagged for parts of speech. If the lengths of nouns and verbs in these corpora allows us to correctly classify Ancient Hebrew and Classical Arabic as VS and the modern varieties as SV, this would suggest that these are salient properties affecting how word order has developed diachronically. Given the textual differences between the taggedPBC (New Testament texts) and the test corpora (Biblical, Quranic, modern, and informal texts), such a finding would be quite striking.
In order to conduct this test, the data for Ancient Hebrew (ISO 639-3: hbo), Modern Hebrew (ISO 639-3: heb), Classical Arabic (cla) [^18] Classical Arabic does not have an assigned ISO 639-3 code or Glottocode. Accordingly, I use the abbreviation cla here, as there is no current language assigned to this code in the ISO 639-3 tables. and Egyptian Arabic (ISO 639-3: arz) were processed and the respective word classes were counted and measured as per the method outlined for Study 1 (each corpus > 400 sentences, see Table 3).
@@
Five features related to word length were used to train a Gaussian Naive Bayes (GNB) classifier, namely the four features described in §2.1.4 above as well as a categorical feature: whether a language’s nouns are longer than its verbs. Using these five features (nounlen, verblen, nlen-vlen, nlen/vlen, nlonger), a GNB classifier was trained on the original combined word order dataset from the taggedPBC containing hand-coded classification of 1,150 languages (Modern Hebrew and Modern Arabic were removed). [^19] As noted previously, since we are concerned here with classification, the collinearity of relevant features is not an issue. The trained classifier was then used to predict word order for our pairs of languages: Ancient Hebrew and Modern Hebrew on the one hand, Classical Arabic and Egyptian Arabic on the other. Ancient Hebrew was classified as VS while Modern Hebrew was classified as SV by the trained classifier. Classical Arabic was classified as VS while Egyptian Arabic was classified as SV.
To double check this finding, I sourced data from another historical language, Old Irish, as represented in the Codex Paulinus Wirziburgensis, whereby Latin text of St. Paul’s epistles was annotated by Irish scholars during the middle of the eighth century. The dataset (Table 4; Doyle 2018) contains over 3,600 Irish glosses/sentences. As an Indo-European language, Old Irish is widely surmised to have changed its word order to verb-initial from an earlier state (c.f. Watkins 1963; Hickey 2002), so this check allows us to verify whether the same properties hold in relation to lengths of nouns and verbs. I then removed Old Irish from the training dataset, retrained the classifier, and predicted on the features of the Würzburg Glosses. Indeed, the classifier accurately classified Old Irish as VS, based on the data extracted from the Würzburg Glosses.
@@
Study 2 thus indicates that the frequency-weighted relative lengths of nouns and verbs can predict word order from corpora unrelated to the type of corpora contained in the taggedPBC, with various implications for our understanding of language processing and language change. This is a notable finding, and I am unaware of any other research that has been able to predict word order change based on measurement of corpus-based features. However, it is possible that this result is conditioned to some degree by the fact that the majority of the test cases are Semitic languages spoken in the same geographical area, so we now turn to an investigation of whether genealogical descent or location provides a better explanation for the correlation of noun and verb lengths with word order.
@@ 2.3 Study 3: Testing variance predicted by word length
The preceding studies have shown, first, that features related to word length differentiate between intransitive word orders crosslinguistically (Study 1) and, second, that such features can predict intransitive word order from corpora for both historical and modern languages (Study 2). While Study 1 controlled for the effect of language family and macroarea on word length features, Study 3 asks a slightly different question. That is: of the factors known to explain word order patterns in languages (family, area, word length), which is more explanatory?
@@
Dunn et al. (2011) showed that descent from a common ancestor was a better predictor of many grammatical patterns than other explanations. Although Jäger & Wahle (2021) counter this claim with additional data, we can also use the current dataset to see if descent is more explanatory of word order for a given language than the lengths of nouns and verbs. Regarding language area, Guzmán Naranjo & Becker (2022) show that areal affects are as important as family membership in predicting linguistic phenomena, and that the results in Dunn et al. can be just as easily predicted by language area. This is also a claim that we can test with the taggedPBC dataset.
@@
To implement the comparison and identify the relative importance of known predictors we can use hierarchical linear regression (HLR). This approach allows researchers to observe the relative impact of particular features that are theoretically relevant to the question at hand by seeing how they influence the performance of successive models. In the case of word order, we can train a base regression model with a feature known to be relevant to word order (the N1 ratio), and then subsequent models that each incorporate additional features of interest (“language area/location”, “family membership”, “noun/verb lengths”). We can then compare the models to see whether the additional features increase the variance explained in the data via the F-value: an increase in F-value (change) and the P-value of this change indicates more/additional variance explained by the particular feature added in a model. Importantly, because we are not concerned with the significance of included features, each individual model can contain multiple (potentially collinear) features related to a given theoretical construct, as what is being measured is simply degree and significance of the change in fit of each model relative to the previous one.
This approach also means we can run multiple analyses with differing sets of data in order to see how this affects the fit of the models. It may be, for example, that using a smaller sample gives a different result than a larger or more diverse one. Here, the more languages we have for a language family or linguistic area, the higher the likelihood that any relationship identified is not due to chance. One caveat of this approach is that when the dependent variable is categorical, an HLR requires that it be transformed to a single distinction. Thus in the ensuing observations I remove languages with “free” word order from consideration, as this leaves us with a binary distinction in word order (SV, VS) compatible with the HLR approach.
@@
Accordingly, for all languages in the taggedPBC I identified family membership and macroarea as noted in §2.1.4. Then, after removing the 63 languages with “free” word order, I filtered the data on four conditions: a) languages found in Dunn et al. that are also present in the taggedPBC (120 languages), b) languages found in the taggedPBC that are in the four families investigated by Dunn et al. (662 languages), c) languages in families with more than 75 members in the taggedPBC (1,182 languages), and d) languages in families with 2 or more members in the taggedPBC (1,805 languages). For each of these conditions, I ran an HLR, adding features of interest progressively to see how they impacted the fit of the successive models and whether particular features account for additional variance.
@@
The base model (Model 1) uses the N1 ratio as the primary feature for classification. This is because we know from comparison with expert determinations of word order (see Ring 2026) that the N1 ratio significantly correlates with the classification of languages as SV or VS. Due to the strong relationship observed between language area and typological properties of language (see Guzmán Naranjo & Becker 2022), Model 2 then includes macroarea identifiers for each language as features. Model 3 additionally incorporates descent (language family membership) as a feature (per Dunn et al. 2011), and Model 4 includes the average lengths of nouns and verbs (per the relationship identified in this paper). [^20] In these models, continuous features (N1 ratio, N/V len) were scaled using the MinMaxScaler from scikit-learn, and categorical features (family, macroarea) were converted to numeric values. If descent is more explanatory than the lengths of nouns/verbs, this should be clear from the differences between the respective models for each condition. Results are displayed in Table 5.
@@
1 (N1 ratio) 120 45.28 6.5e-10
2 (Lg area) 120 46.63 1.28e-15 34.95 3.41e-08 **
3 (Family) 120 30.86 9.65e-15 0.07 0.79
4 (N/V len) 120 19.57 4.92e-14 1.91 0.15
@@
1 (N1 ratio) 662 190.0 3.61e-38
2 (Lg area) 662 182.17 1.05e-63 135.59 1.25e-28 **
3 (Family) 662 122.02 7.73e-63 1.47 0.23
4 (N/V len) 662 82.34 4.44e-67 15.02 4.18e-07 **
@@
1 (N1 ratio) 1182 378.83 2.14e-73
2 (Lg area) 1182 327.01 1.05e-113 208.55 1.2e-43 **
3 (Family) 1182 218.01 1.79e-112 0.37 0.54
4 (N/V len) 1182 132.79 1.16e-111 3.55 0.03 *
@@
1 (N1 ratio) 1805 535.42 6.17e-104
2 (Lg area) 1805 327.44 4.99e-122 92.34 2.35e-21 **
3 (Family) 1805 218.53 6.92e-121 0.78 0.38
4 (N/V len) 1805 136.84 3.62e-123 10.76 2.27e-05 **
@@
From these results we can see that, as expected, the N1 ratio significantly predicts word order under all conditions. Additionally, language area is a significant predictor of word order that adds to the strength of the classification. Language family (descent) does not add additional predictive power to our model under most conditions (over language area), whereas the length of nouns and verbs does. This analysis indicates that the length of nouns and verbs in a language accounts for more variance and at the same time is more significant than that language’s family membership in determining its word order classification, under all test conditions.
When we remove language area from the linear regression, [^21] See additional statistical results at the linked repositories (Fn 2; Fn 5). we find that family becomes a significant predictor of word order for a smaller sample of languages, but that under these conditions the length of nouns and verbs still accounts for additional variance and is also highly significant. Under conditions (c) and (d), with larger and more diverses samples, the model incorporating language family provides little explanation and is non-significant, while the subsequent model incorporating length of nouns and verbs explains additional variance that is highly significant.
This largely aligns with Jäger and Wahl’s (2021) finding regarding the (lower) correlation between descent and linguistic structures in comparison to other explanations, with the caveat that Jäger and Wahl included language isolates, which is likely to greatly diminish the affect of descent. If a large number of languages in a sample are not related to others, clearly descent can have only a minimal impact on word order. By including only language families with two or more members in the taggedPBC (which excludes 74 languages classified as having either SV/VS order), I show that descent does not explain much variance above and beyond language location/area, in contrast with noun and verb lengths. The test where language area was removed altogether from the HLR models also suggests that language family/descent is highly correlated with linguistic area, which conforms with expectations related to human history and migration. Additionally, the amount of variance explained seems to depend on the number of languages (and language families) in the sample, highlighting the need for more data for languages from under-represented language families.
@@ 3 Discussion
The preceding investigations have shown that the average length of nouns and verbs, weighted by their frequency in a corpus, correlates with intransitive word order crosslinguistically. This correlation is strong enough that such word length features can predict word order for both historical and modern languages, and is a stronger signal for word order than either language family or language area. These findings have various implications for our understanding of language processing and how languages change under an agentic evolutionary model. In particular, the effect of frequency and its relation to processing suggests some possibilities for reconceptualizing a theory of language behavior.
The fact that a clear relationship between word order and word length is not observed unless frequency is taken into account indicates that the availability of corpus data is required to investigate questions related to crosslinguistic grammatical patterns and how they might develop. Here, the ability to group classes of words for observation allows for more fine-grained analyses regarding such patterns. For example, when we group word classes into arguments (prototypical agents/actors as represented by nouns, pronouns, proper nouns) and predicates (prototypical actions/states as represented by verbs, auxiliaries) there is a different realization of length than if we solely consider nouns and verbs. In the former condition all arguments are shorter than verbs crosslinguistically, while in the latter condition the differences in length emerge.
That arguments overall are shorter in languages of the taggedPBC, across all word orders, suggests that there are certain (non-Noun) arguments affecting these observations. While this has yet to be investigated in detail, the effect of PRON seems quite likely, whereby a tendency for pronouns to be shorter crosslinguistically, along with a preference for pronouns as arguments, may impact the observation that arguments are shorter overall than predicates. This pattern aligns with the processing claim that “heavier” words (or word classes in this case) occur later in a sentence, partly validating Hypothesis 1 of study 1. The counter-argument, also from the data, is that nouns are longer than verbs in SV languages and shorter than verbs in VS languages when frequency is considered, validating Hypothesis 2 of study 1. This suggests that speakers prefer to place heavier (more “contentful”, or perhaps “novel”) constituents before lighter ones crosslinguistically in many cases.
@@
These two pressures can be (and, in fact, have been) considered as two different kinds of “efficiency”: production efficiency vs communicative efficiency. Here production efficiency is related to length of words (as per Wasow 2002, 2022; Hawkins 2014) - speakers use shorter words whenever they can to efficiently produce utterances (see also Mahowald et al. 2013). The counter-pressure, communicative efficiency, is related to making a speaker’s utterances more readily accessible/interpretable by recipients (as per Levshina & Moran 2021; Levshina 2023) - speakers prioritize words that contain more meaningful, informative content in order to aid communication. Once context has been established, these longer words can be replaced by more efficient (shorter) terms such as demonstratives or pronouns. [^22] For example, there is a common crosslinguistic pattern of using a noun first to establish reference, then switching to pronouns, i.e. "that tree, it is big, it will fall…" See also discussion of the (lack of) encoding of subjects crosslinguistically in Berdicevskis et al. (2020).